Write a python program to design a function with a parameter called A (which is between 1 and 10) and prints the multiplication table for 1 to A. • Note: Do not need to draw the horizontal and vertical lines. • Example for A = 10
Answer #1

Code:

def A(n):
    for i in range(1,n+1):
        for j in range(1,11):
            v = i * j#calculating the product
            print('{} * {} = {}'.format(i,j,v))
        print()#newline

A(5)

output for parameter value = 5:

Pratyushs-MacBook-Pro:python pthapli$ python3 main.py 1 * 1 = 1 1 * 2 = 2 1 * 3 = 3 1 * 4 = 4 1 * 5 = 5 1 * 6 = 6 1 * 7 = 7 1

Code screenshot: